Find the number halfway between the given numbers.
1.) 7/11 and 3/4

Found 2 solutions by Alan3354, macston:
Answer by Alan3354(69443)

Find the number halfway between the given numbers.
1.) 7/11 and 3/4
---------------
Find a common denominator to use.
Then find the average of the 2 numerators over the common DEN.

Answer by macston(5194)

To find the halfway point, find a common denominator, add the numbers, and divide by two.
(7/11)(4/4)=28/44
(3/4)(11/11)=33/44)
Find the sum:
(28/44)+(33/44)=61/44
Divide by 2:
(61/44)(1/2)=61/88 ANSWER: The number halfway between 7/11 and 3/4 is 61/88
CHECK:
(28/44)(2/2)=56/88
61/88-56/88=5/88
(33/44)(2/2)=66/88
66/88-61/88=5/88
61/88 is 5/88 from 7/11 and 5/88 from 3/4, so it is the halfway point.